# SteerMusic: Enhanced Musical Consistency for Zero-shot Text-Guided and Personalized Music Editing

### v1

@@ 4.2. SteerMusic+: Personalized Music Editing
Personalized Diffusion Model (PDM) serves a foundational part in SteerMusic+. Since training a PDM has been extensively studied in (Plitsis et al., 2024; Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023), we assume the availability of a pretrained text-to-music PDM in SteerMusic+, as SteerMusic+ is a plug-in pipeline compatible with existing pretrained PDM. The text-to-music PDM $\epsilon_{\phi^{\prime}}$ captures the user-defined musical concept $S$ by fine-tuning an existing pretrained DPM $\epsilon_{\phi}$ under a small set of reference music $\mathcal{D}^{\text{ref}}=\{(x^{\text{ref}},y^{\text{ref}})_{n}\}^{N}_{n=1}$, where $N$ can be as few as $1$ (Plitsis et al., 2024). The fine-tuning is achieved via optimizing the objective
(4) $\phi^{\prime}\in\text{arg min}_{\hat{\phi}}\mathbb{E}_{(x^{\text{ref}},y^{\text{ref}})\sim\mathcal{D}^{\text{ref}}}\mathcal{L}_{\text{DPM}}(x^{\text{ref}},y^{\text{ref}};\hat{\phi})$
where $\hat{\phi}$ is initialized with a pretrained DPM weights $\phi$. As illustrated in Figure 4 (a), the prompt $y^{\text{ref}}$ takes form of “a recording of a [$S$]”, where the placeholder [$S$] corresponds to a defined new concept word embedding. During inference, the PDM can generate music with the newly learned concept (e.g., “A disco song with a [$S$]”). The dataset $\mathcal{D}^{\text{ref}}$ consists of reference audios that encapsulate the concept users aim to extract. According to (Plitsis et al., 2024), the concept represents a musical style, which can be either an instance of instrument sounds or a specific genre that cannot be yielded even with the most detailed textual description. For instance, if the objective is to capture the user’s guitar playing style, the reference audio should feature performances on the user’s guitar playing. Conversely, if the goal is to capture the concept of jazz, the reference audio should consist of recordings that exemplify the same jazz style.
Although PDM can generate new music compositions with new concepts captured from the reference data set $\mathcal{D}^{\text{ref}}$, it cannot modify existing music using $y^{\text{tgt}}$ containing the new concepts directly. Next, we introduce the key components of SteerMusic+ that enable the PDM to perform personalized music editing.
Personalized Delta Score (PDS) is an extension of Eq. 3 to enable personalized music editing. We define the PDS loss as $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,\\ x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t)\|^{2}_{2}]$. By omitting the UNet Jacobian, the gradient over $\theta$ is given by
(5) $\nabla_{\theta}\mathcal{L}_{\text{PDS}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$
where $x_{t}$ and $x_{t}^{\text{src}}$ shares the same sampled noise $\epsilon$ at time step $t$.
This modified delta score can be decomposed to two components: The score of $\epsilon_{\phi^{\prime}}$ provides a desired direction to guide the editing to match the target prompt with concept $[S]$. The score of $\epsilon_{\phi}$ helps to reduce noisy direction of unintended modification areas. The delta score between the PDM $\epsilon_{\phi^{\prime}}$ and the DPM $\epsilon_{\phi}$ may not yield an effective direction toward $y^{\text{tgt}}$, as $\epsilon_{\phi^{\prime}}$ shifted to the reference distribution $\mathcal{D}^{\text{ref}}$. We introduce an additional component to compensate for the distribution shift induced by the score of $\epsilon_{\phi^{\prime}}$.
Distribution Shift Regularization. To bridge the distribution gap between $\epsilon_{\phi^{\prime}}$ and $\epsilon_{\phi}$, we introduce a regularization term to further regularize the edited score to the personalized diffusion model $\epsilon_{\phi^{\prime}}$. We wish to minimize the distribution shift between two diffusion models by adding a constraint as
(6) $\min_{\theta}$ $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}}),$
subject to $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})-\zeta\leq 0.$
where $\zeta$ is a small amount of constant, the regularization term is $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)\|^{2}_{2}]$. Same as above, the gradient with respect to $\theta$ is given by
(7) $\nabla_{\theta}\mathcal{L}_{\text{shift}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t))\frac{\partial x}{\partial\theta}]$
By integrating the constraint, the overall gradient through $\theta$ is
(8) ${}\nabla_{\theta}\mathcal{L}_{\text{PDS-O}}(\phi^{\prime},\phi,x,y,x^{\text{src}},y^{\text{src}})=\underbrace{\nabla_{\theta}\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})}_{\text{`Delta score point to edit direction'}}\\ +\lambda\underbrace{\nabla_{\theta}\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})}_{\text{`Delta score regularizes distribution shift'}}$
where $\lambda$ is a predefined constant that adjusts the strength of the regularization. We will discuss the effect of $\lambda$ in our ablation study.
Table 1. Model comparison on zero-shot text-guided music editing task.
@@
DDIM (Song et al., [n. d.]) 0.477 1.022 0.330 5.377 0.264 1.37 1.91
SDEdit (Meng et al., 2021) 0.638 1.274 0.169 6.208 0.218 0.92 1.68
MusicMagus (Zhang et al., 2024b) 0.593 1.698 0.338 5.243 0.238 2.11 1.57
ZETA (Manor and Michaeli, 2024) 0.509 1.021 0.293 5.458 0.252 1.22 1.60
SteerMusic (Ours) 0.278 0.381 0.480 3.772 0.259 2.92 2.50
@@
Personalized Contrastive (PCon) Loss. Eq. 8 formulates a regularized editing direction, where the regularized delta score encourages alignment with the target prompt while mitigating the distribution shift. It does not explicitly enforce fidelity to the concept, which may lead to suboptimal editing quality. To further enhance the edit fidelity, we incorporate a PCon loss between temporal features, which is modified from a patch-wise contrastive loss (Nam et al., 2024). PCon loss extracts intermediate features $h^{\text{src}}_{l}$ and $h_{l}$ that passed through the residual block and self-attention block from $\epsilon_{\phi}$ conditioned on $y^{\text{src}}$ and $\epsilon_{\phi^{\prime}}$ conditioned on $y^{\text{tgt}}$ at $l$-th self-attention layer, respectively. The features are then reshaped into size $\mathbb{R}^{T_{l}\times F_{l}\times C_{l}}$, where $T_{l},F_{l}$, and $C_{l}$ represent the size of the temporal, spatial, and channel dimension in the $l$-th layer, respectively. The patch corresponding to the temporal location on the feature map $h^{\text{src}}_{l}$ is designated as ‘positive’, and vice versa. The PCon loss is defined as
(9) $\mathcal{L}_{\text{PCon}}(x,x^{\text{src}})=\mathbb{E}_{h}[\sum_{l}\sum_{t^{\prime}}\ell(h_{l}^{t^{\prime}},h_{l}^{\text{src},t^{\prime}},h_{l}^{\text{src},T_{l}\backslash t^{\prime}})]$
(10) $\ell(h,h^{+},h^{-})=-\text{log}(\frac{\text{exp}(h\cdot h^{+}/\tau)}{\text{exp}(h\cdot h^{+}/\tau)+\text{exp}(h\cdot h^{-}/\tau)})$
@@
where $t^{\prime}\in\{1,...,T_{l}\}$ represents temporal location query patch, the positive patch as $h_{l}^{src,t^{\prime}}$ while the other patches as $h_{l}^{src,T_{l}\backslash t^{\prime}}$. $\text{exp}(h\cdot h^{+}/\tau)$ is positive sample with the same temporal location, $\text{exp}(h\cdot h^{-}/\tau)$ is negative sample with mismatched temporal location in the self-attention features, $\tau$ is a temperature parameter as $\tau>0$.
@@
The gradient of $\mathcal{L}_{\text{PCon}}(x,x^{\text{src}})$ will propagate to the hidden state of self-attention layers $h$ in the personalized diffusion $\epsilon_{\phi^{\prime}}$. Given that the personalized diffusion $\epsilon_{\phi^{\prime}}$ has a distribution shift over the reference dataset $\mathcal{D}^{\text{ref}}$, the $\mathcal{L}_{\text{PCon}}$ explicitly encourages feature similarity at the frequency domain in self-attentions, particularly for attributes that distinguish the target concept. This reinforcement leads the model to prioritize concept consistency over strict temporal alignment with the source music. Consequently, $\mathcal{L}_{\text{PCon}}$ amplifies the distinctive characteristics of the target concept of $\epsilon_{\phi^{\prime}}$ in SteerMusic+, ensuring that the edited music maintains stronger fidelity to the desired style while allowing structural variations.
5. Experiment
5.1. Experimental Setup
To evaluate our methods, we use a pretrained AudioLDM2 (Liu et al., 2024b) as the foundational model. In the text-guided music editing experiment, we set the optimization iteration as 400 with a 30 guidance scale (Ho and Salimans, [n. d.]) in the diffusion model on Eq. 3 in SteerMusic. In personalized music editing experiment, we follow the setting of the official repository of DreamSound [^1] https://github.com/zelaki/DreamSound and fine-tune the DreamSound model 100 steps with $1e^{-5}$ learning rate. We set 400 optimization steps with 15 guidance scale and 0.05 $\lambda$ value in Eq. 8 of SteerMusic+. All experiments were performed on a single NVIDIA H100 GPU.
Figure 5. A top CQT feature comparison between source and text-guided edited music for different methods. SteerMusic (ours) produces edited results have similar top CQT compare to other methods, indicating SteerMusic successfully preserves source music melody.
Figure 6. Radar chart of model comparison for each text-guided music editing subtask on SteerMusic.
5.2. Evaluation Metrics
@@
We evaluate music editing objectively based on two aspects: musical consistency (content preservation before and after editing) and editing fidelity. We follow (Manor and Michaeli, 2024) and calculate the following objective metrics for measurement. To evaluate musical consistency, we use:
@@
Fréchet Audio Distance (FAD) (Kilgour et al., 2019) measures the distributional difference between source and edited music (lower the better). We calculate FAD based on both VGGish (Hershey et al., 2017) and clap-laion-music (Wu et al., 2023) embeddings. LPAPS (Iashin and Rahtu, 2021), an audio version of LPIPS (Zhang et al., 2018) which quantifies the consistency of the edited audio relative to the source audio (lower the better). Top-1 Constant-Q Transform Pearson Correlation Coefficient (CQT-1 PCC) measures melody consistency between the source and edited music (higher the better). The CQT-1 (Brown, 1991) extracts main melody of the music audio and has been shown to outperform traditional chroma-based features in representing melodic characteristics (Hou et al., 2025). While existing metrics such as FAD and LPAPS provide insight into audio quality and perceptual similarity, they fall short in comprehensively capturing melodic structure. Although existing deep learning models (Cwitkowitz et al., 2024; Bittner et al., 2022; Gardner et al., 2022; Chang et al., 2024; Mancusi et al., 2025) effectively extract melodies from real music, they are unreliable for synthesized audio. To address the limitation, we introduce CQT-1 PCC as a supplementary objective metric, specifically designed to quantify melodic consistency in generative music editing. This metric enables a more targeted evaluation of whether the core melodic structure of the source audio is retained in the edited result. More details of the metrics are provided in the Supplementary Material.
@@
CLAP Score (Wu et al., 2023; Chen et al., 2022) measures alignment between edited music and the target prompt in text-guided music editing (higher is better). CDPAM (Manocha et al., 2021), a perceptual audio metric leveraging deep learning representations, measures perceptual distance between audios such as music and speech (Jacobellis et al., 2024; Hai et al., 2024; Niu et al., 2024; Gui et al., 2024; Chung et al., 2025). We use CDPAM to evaluate audio perceptual similarity between reference music and the edited result in personalized music editing (lower the better).
Subjective evaluation: To further validate our methods, we designed a Mean Opinion Score (MOS) study to evaluate both target editing fidelity (MOS-T) and source content preservation (MOS-P) subjectively by asking participants to rate results from 1-Bad to 5-Excellent (ITU-T RECOMMENDATION, 1999; Sector, 1996) for randomly selected edited samples.222 This study has been reviewed by the internal privacy committee to ensure compliance with local regulations for ethical conduct of research with human participants and ACM’s Code of Ethic. More details are provided in the Supplementary Material.
5.3. Zero-shot Text-guided Music Editing
In this part, we evaluate our SteerMusic method proposed in Sec. 4.1 on zero-shot text-guided music editing task.
Dataset: We use the ZoME-Bench (Liu et al., 2024a) dataset [^3] https://huggingface.co/datasets/liuhuadai/ZoME-Bench to evaluate our method, which contains 1000 10-second audio samples obtained from MusicCaps (Agostinelli et al., 2023), with corresponding source and target prompts for the editing. ZoME-Bench has 9 subdatasets that target different music editing tasks. Since SteerMusic primarily focuses on enhancing musical consistency for music style editing, we select specific subsets from the dataset that involve instrument changes, genre changes, mood changes, and background modifications. These subsets contain 131, 134, 100, and 95 music clips, respectively.
Baseline: We directly compare SteerMusic with other plug-in zero-shot text-guided music editing methods, SDEdit (Wang et al., 2023a; Meng et al., 2021), DDIM (Song et al., [n. d.]), ZETA (Manor and Michaeli, 2024) and MusicMagus (Zhang et al., 2024b), which based on the same pretrained AudioLDM2 (Liu et al., 2024b). We do not compare with MEDIC (Liu et al., 2024a) and MelodyFlow (Le Lan et al., [n. d.]) as their source code has not been released. We also exclude AudioEditor (Jia et al., 2025) and AudioMorphX (Liang et al., 2024), which are primarily designed for sound editing rather than music editing.
Table 2. Model comparison on personalized music editing task.
@@ 5.3. Zero-shot Text-guided Music Editing
Textual inv. (Plitsis et al., 2024) 0.789 1.871 0.216 5.083 0.713 1.64 1.63
DreamSound (Plitsis et al., 2024) 0.902 3.175 0.292 5.082 0.609 1.42 1.81
SteerMusic+ (Ours) 0.362 0.558 0.399 4.125 0.593 3.07 2.47
Figure 7. User preference study for SteerMusic. We report the percentage of users that preferred SteerMusic (ours) over ZETA (Manor and Michaeli, 2024) and MusicMaugs (Zhang et al., 2024b) in this user study.
Figure 8. A visualization of edited results between SteerMusic+ and baselines in personalized music editing. SteerMusic+ preserves instruction-irrelevant musical content on the source music.
Experimental results: Table 1 compares SteerMusic with other zero-shot text-guided music editing baselines across various music style transfer tasks. Our method demonstrates notable improvements in musical consistency, as evidenced by higher CQT-1 PCC, lower LPAPS, and lower FAD scores. DDIM achieves the best CLAP score (0.264), it has a much lower CQT-1 PCC and a significantly higher LPAPS. This suggests that while DDIM achieves higher alignment with the target style in terms of text-audio correspondence, it directly conditions on the target prompt during the denoising process without applying constraints to preserve the original music information. In contrast, SteerMusic balances both source music consistency and target style fidelity, ensuring that the resulting music remains faithful to the original music while effectively incorporating the target style changes. Moreover, subjective evaluation scores (MOS-P and MOS-T) further support the superiority of SteerMusic in both content preservation and edit fidelity, aligning well with the objective metrics. Figure 5 presents a visualized comparison of top-1 CQT features for an instrument editing sample. SteerMusic generates results with CQT features that closely resemble those of the original music, suggesting that it effectively preserves the original music melody consistency. A more detailed comparison across individual subtasks is in Figure 6.444See our Supplementary Material for extra comparison with SDS (Poole et al., 2022) and CDS (Nam et al., 2024).
@@
User preference study: We evaluate SteerMusic by a user preference study following the design in (Manor and Michaeli, 2024)2. To reduce cognitive load and improve reliability, we compare SteerMusic with the top-performing baselines, MusicMagus and ZETA. In this study, users were asked to answer a sequence of 20 questions, each question contains original music, an editing instruction, and two edited results. Users were instructed to select the edited result that matches the instruction better while preserving the main original music content. We collected 25 full responses, which the participants having a minimum of 1 and average of 5 years of music training. Our method was clearly preferred over all competing methods (see Figure 7).
5.4. Personalized Music Editing
@@
In this part, we evaluate our SteerMusic+ method proposed in Sec. 4.2 on personalized music editing task.
@@
Dataset: We selected a representative subset of eight musical concepts from the 32 concepts provided by (Plitsis et al., 2024), for both common and uncommon cases, which include four instrument concepts (i.e., Guitar, Bouzouki, Ocarina, Sitar) and four genre concepts (i.e., Morricone, Reggae, Hiphop, Sarabande). Each musical concept includes an instruction of concept placeholder and five 10-second reference audio clips sourced from YouTube and FreeSound. For the source music on the personalized editing, we follow Sec. 5.3 and use the “change instruments” and “change genre” subsets in ZoME-banch (Liu et al., 2024a) to perform personalized instrument transfer and personalized genre transfer, which contains 131 and 134 audio clips respectively. We manipulate the concept token by adapting the original target prompt provided by ZoME-banch for each audio sample from the original word replacement to the musical concept.

### v3

@@ Introduction
Personalized Diffusion Model (PDM) serves a foundational part in SteerMusic+. Since training a PDM has been extensively studied in plitsis2024investigating; ruiz2023dreambooth; gal2022image; kumari2023multi, we assume the availability of a pretrained text-to-music PDM in SteerMusic+, as SteerMusic+ is a plug-in pipeline compatible with existing PDMs. The text-to-music PDM $\epsilon_{\phi^{\prime}}$ captures the user-defined musical concept $S$ by fine-tuning a pretrained DPM $\epsilon_{\phi}$ under a small set of reference music $\mathcal{D}^{\text{ref}}=\{(x^{\text{ref}},y^{\text{ref}})_{n}\}^{N}_{n=1}$, which $N$ can be as few as $1$ (plitsis2024investigating). The fine-tuning is achieved via optimizing the objective
@@
where $\hat{\phi}$ is initialized with a pretrained DPM weights $\phi$. As illustrated in Fig. 4 (a), the prompt $y^{\text{ref}}$ takes form of “a recording of a [SS]”, where the placeholder [$S$] corresponds to a defined new concept word embedding. During inference, the PDM can generate music with the newly learned concept (e.g., “A disco song with a [SS]”). The dataset $\mathcal{D}^{\text{ref}}$ consists of reference audio clips that encapsulate the concept users aim to extract. According to plitsis2024investigating, the concept represents a musical style, which can be either an instance of instrument sounds or a specific genre that cannot be yielded even with the most detailed textual description. For instance, if the objective is to capture the user’s guitar playing style, the reference audio should feature performances on the user’s guitar playing. Conversely, if the goal is to capture the concept of jazz, the reference audio should consist of recordings that exemplify the same jazz style.
While the PDM can generate new music based on concepts from $\mathcal{D}^{\text{ref}}$, it cannot directly edit existing music using $y^{\text{tgt}}$ that contains the concepts. Next, we introduce key components of SteerMusic+ that enable personalized editing with PDMs.
Personalized Delta Score (PDS) is an extension of Eq. 1 to enable personalized music editing. We define the PDS loss as $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x=g(\theta),y^{\text{tgt}},x^{\text{src}},y^{\text{src}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t)\|^{2}_{2}]$. By omitting the UNet Jacobian, the gradient over $\theta$ is given by $\nabla_{\theta}\mathcal{L}_{\text{PDS}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$, where $x_{t}$ and $x_{t}^{\text{src}}$ share the same sampled noise $\epsilon$ at the time step $t$. This modified delta score can be decomposed into two components: the score of $\epsilon_{\phi^{\prime}}$ provides the desired direction to guide the editing to match the target prompt with the concept $[S]$. The score of $\epsilon_{\phi}$ reduces the noisy direction of unintended modification areas. The delta score between the PDM $\epsilon_{\phi^{\prime}}$ and the DPM $\epsilon_{\phi}$ may not produce an effective direction toward $y^{\text{tgt}}$, as $\epsilon_{\phi^{\prime}}$ shifted to the reference distribution $\mathcal{D}^{\text{ref}}$. We introduce an additional component to compensate for the distribution shift induced by the score of $\epsilon_{\phi^{\prime}}$.
Distribution Shift Regularization. To bridge the distribution gap between $\epsilon_{\phi^{\prime}}$ and $\epsilon_{\phi}$, we introduce a regularization term to regularize the edited score to the personalized diffusion model $\epsilon_{\phi^{\prime}}$. We wish to minimize the distribution shift between two diffusion models by adding a constraint as
$\min_{\theta}$ $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x=g(\theta),y^{\text{tgt}},x^{\text{src}},y^{\text{src}}),$ (3)
subject to $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x=g(\theta),y^{\text{tgt}})-\zeta\leq 0.$
where $\zeta$ is a small amount of constant; the regularization term is $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x=g(\theta),y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)\|^{2}_{2}]$. The gradient in respect to $\theta$ is given by $\nabla_{\theta}\mathcal{L}_{\text{shift}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t))\frac{\partial x}{\partial\theta}]$. Therefore, the overall gradient through $\theta$ is
[…]
where $\lambda$ is a constant that adjusts regularization strength.
@@
Personalized Contrastive (PCon) Loss. Eq. 4 formulates a regularized reference guided editing direction, where the regularized delta score encourages alignment with the target prompt while mitigating the distribution shift. However, it does not explicitly enforce the fidelity to the concept, which may lead to suboptimal editing quality. To further enhance the fidelity of the edit, we incorporate a PCon loss between temporal features, which is modified from a patch-wise contrastive loss (nam2024contrastive). PCon loss extracts intermediate features $h^{\text{src}}_{l}$ and $h_{l}$ that pass through the residual block and the self-attention block from $\epsilon_{\phi}$ conditioned on $y^{\text{src}}$ and $\epsilon_{\phi^{\prime}}$ conditioned on $y^{\text{tgt}}$ at the $l$-th self-attention layer, respectively. The features are then reshaped to size $\mathbb{R}^{T_{l}\times F_{l}\times C_{l}}$, where $T_{l},F_{l}$, and $C_{l}$ represent the size of the temporal, spatial, and channel dimensions in the $l$-th layer, respectively. The patch corresponding to the temporal location on the feature map $h^{\text{src}}_{l}$ is designated as ‘positive’, and vice versa. The PCon loss is defined as
[…]
$\ell(h,h^{+},h^{-})=-\text{log}(\frac{\text{exp}(h\cdot h^{+}/\tau)}{\text{exp}(h\cdot h^{+}/\tau)+\text{exp}(h\cdot h^{-}/\tau)})$
@@
where $t^{\prime}\in\{1,...,T_{l}\}$ represents the temporal location query patch, the positive patch as $h_{l}^{src,t^{\prime}}$ while the other patches as $h_{l}^{src,T_{l}\backslash t^{\prime}}$. $\text{exp}(h\cdot h^{+}/\tau)$ is a a positive sample with the same temporal location, $\text{exp}(h\cdot h^{-}/\tau)$ is a negative sample with a mismatched temporal location in the self-attention features, $\tau$ is a temperature parameter as $\tau>0$.
@@
The gradient of $\mathcal{L}_{\text{PCon}}(x,x^{\text{src}})$ will propagate to the hidden state of self-attention layers $h$ in personalized diffusion $\epsilon_{\phi^{\prime}}$. Given that the personalized diffusion $\epsilon_{\phi^{\prime}}$ has a distribution shift over the reference dataset $\mathcal{D}^{\text{ref}}$, the $\mathcal{L}_{\text{PCon}}$ explicitly encourages feature similarity at the frequency domain in self-attentions, particularly for attributes that distinguish the target concept. This reinforcement leads the model to prioritize concept consistency over strict temporal alignment with the source music. Consequently, $\mathcal{L}_{\text{PCon}}$ amplifies the distinctive characteristics of the target concept of $\epsilon_{\phi^{\prime}}$ in SteerMusic+, ensuring that the edited music maintains stronger fidelity to the desired style while allowing structural variations.
Experiments
[…]
DDIM 0.477 4.713 0.330 5.377 0.264 1.37 1.91
SDEdit 0.638 6.749 0.169 6.208 0.218 0.92 1.68
MusicMagus 0.593 7.631 0.338 5.243 0.238 2.11 1.57
ZETA 0.509 3.380 0.293 5.458 0.252 1.22 1.60
SteerMusic 0.278 2.426 0.480 3.772 0.259 2.92 2.50
Table 1: Model comparison on zero-shot text-guided music editing task using the ZoME-Bench dataset.
Evaluation Metrics
@@
We evaluate music editing objectively based on two aspects: musical consistency (content preservation before and after editing) and editing fidelity. We follow manor2024zero and calculate the following objective metrics for measurement. To evaluate musical consistency, we use:
[…]
  • LPAPS (iashin2021taming), an audio version of LPIPS (zhang2018unreasonable), which quantifies the consistency of the edited audio relative to the source audio (lower the better).
@@
  • Top-1 Constant-Q Transform Pearson Correlation Coefficient (CQT1-PCC), which measures melody consistency between the source and edited music (higher the better). The CQT1-PCC (brown1991calculation) extracts the main melody of the music audio and has been shown to outperform traditional chroma-based features in representing melodic characteristics (hou2024editing). While existing metrics such as FAD and LPAPS provide insight into audio quality and perceptual similarity, they fall short in comprehensively capturing melodic structure. Moreover, existing transcription models (cwitkowitz2024timbre; bittner2022lightweight; gardner2022mt; chang2024yourmt3+; mancusi2025latent) effectively extract melodies from real music, they are unreliable for synthesized audio. To address these limitations, we introduce CQT1-PCC as a supplementary objective metric, specifically designed to quantify melodic consistency in generative music editing. This metric enables a more targeted evaluation of whether the core melodic structure of the source audio is retained in the edited result.
@@
  • CLAP Score (wu2023large) which measures the alignment between edited music and the target prompt in text-guided music editing (higher is better).
  • CDPAM (manocha2021cdpam), a perceptual audio metric that leverages deep learning representations to measure perceptual distance between audios such as music and speech (jacobellis2024machine; hai2024dpm; gui2024adapting). We use CDPAM to evaluate audio perceptual similarity between reference music and the edited result in personalized music editing (lower the better).
Subjective evaluation. We designed a mean opinion score (MOS) study to evaluate target editing fidelity (MOS-T) and source content preservation (MOS-P) by asking participants to rate the results from 1-Bad to 5-Excellent (international1996methods) for randomly selected edited samples. We provide more experimental details in our extended version.
Zero-shot Text-guided Music Editing
In this part, we evaluate our SteerMusic method on the zero-shot text-guided music editing task.
Dataset. We use the ZoME-Bench dataset (liu2024medic) which includes 1,000 10-second audio samples from MusicCaps (agostinelli2023musiclm), each paired with source and target text prompts. We evaluate our models on four well-defined editing tasks that require modifying a specific aspect of the audio while preserving the original melody: change instrument (131 clips), change genre (134), change mood (100), and change background (95). To assess long-form editing, we use the MusicDelta subset of MedleyDB (musicdelta) with ranging from 20 seconds to 5 minutes, comprising 34 excerpts of varying styles and lengths, with prompts from manor2024zero.
Baseline. We compare SteerMusic with zero-shot text-guided music editing methods plug-in the same pretrained AudioLDM2 (liu2024audioldm), including SDEdit (meng2021sdedit), DDIM (song2020denoising), ZETA (manor2024zero), and MusicMagus (zhang2024musicmagus). To ensure statistical reliability, experiments are conducted using multiple random seeds. We are unable to include MelodyFlow (le2024high) and MEDIC (liu2024medic) due to the lack of source code. We exclude AudioEditor (jia2025audioeditor) and AudioMorphX (liang2024audiomorphix), which are designed for general sound editing rather than music.
Figure 5: User preference for SteerMusic: percentage of users preferring our method over ZETA and MusicMagus.
Figure 6: A visualization of edited results between SteerMusic+ and baselines in personalized music editing. SteerMusic+ preserves instruction-irrelevant musical content on the source music.
@@
Textual Inv. 0.789 9.688 0.216 5.083 0.713 1.64 1.63
DreamSound 0.902 10.686 0.292 5.082 0.609 1.42 1.81
SteerMusic+ 0.362 4.434 0.399 4.125 0.593 3.07 2.47
Table 2: Model comparison on personalized music editing task using the ZoME-Bench dataset.
Experimental results. Tab.1 compares SteerMusic with zero-shot baselines across various style transfer tasks. SteerMusic achieves higher source consistency, shown by improved CQT1-PCC, lower LPAPS, and FAD. While DDIM attains a slightly higher CLAP score (+5e-3), its low CQT1-PCC and high LPAPS indicate poor preservation of source content due to lack of further source consistency constraints during denoising. In contrast, SteerMusic effectively balances source consistency and edit fidelity, fulfilling the core objective of music editing. Furthermore, our method attains the highest MOS scores and yields statistically significant improvements over all baseline models. An ANOVA test shows MOS-P =2.92=2.92 with p-value $7.37\times 10^{-27}$ and MOS-T =2.5=2.5 with p-value $3.24\times 10^{-7}$. These results demonstrate that SteerMusic provides substantially better editing performance in terms of both source-music consistency and editing fidelity as perceived by human listeners.
To assess real-world applicability, we further evaluate our method on MusicDelta dataset in Tab. 3, which consists of varying lengths of music clips. MusicMagus fails in this experiment as it was designed for 5-second editing and doesn’t support long-form music, and hence is excluded. SteerMusic consistently outperforms other baselines in terms of edit fidelity and source consistency, demonstrating its robustness and effectiveness in handling longer and more complex music editing.
@@
User preference study. We evaluate SteerMusic with the user preference study following the design in manor2024zero. To reduce cognitive load and improve reliability, we compare SteerMusic with the top-performing baselines, MusicMagus and ZETA. In this study, users were asked to answer a sequence of 20 questions, each question contains original music, an editing instruction, and two edited results. Users were instructed to select the edited result that better matches the instruction while preserving the main content of the original music. We collected 25 full responses, which the participants having a minimum of 1 and average of 5 years of music training. As shown in Fig. 5, our method was clearly preferred over all competing methods.
Personalized Music Editing
@@
In this part, we evaluate our SteerMusic+ method, which is designed for personalized music editing task.
@@
Dataset. To cover both common and exotic concepts, we selected eight representative musical concepts from the 32 defined in plitsis2024investigating: four instruments style (Guitar, Bouzouki, Ocarina, Sitar) and four genres (Morricone, Reggae, Hiphop, Sarabande). Each concept includes a placeholder instruction and five 10-second reference clips from YouTube and FreeSound. We use the “change instrument” and “change genre” tasks from ZoME-Bench (liu2024medic), replacing the original target prompt with the selected concept token. Due to the lack of detailed instructions in MusicDelta, we use the standardized prompt as “A recording of a [style] song.”, where [style] is either the source style or the target style concept [S].
[…]
DDIM 0.646 3.336 0.245 4.972 - 0.316
ZETA 0.665 3.789 0.296 5.071 - 0.319
SDEdit 0.818 8.757 0.137 5.996 - 0.310
SteerMusic 0.622 2.559 0.351 4.122 - 0.321
DreamSound 0.847 8.972 0.220 5.318 0.583 -
SteerMusic+ 0.666 5.506 0.273 4.574 0.581 -
Table 3: Model comparison of SteerMusic and SteerMusic+ on MusicDelta dataset.